OpenStudy (anonymous):

Simplify √28 (Use sq rt as the square root symbol)

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{28}\] \[\Large \sqrt{4*7}\] \[\Large \sqrt{4}*\sqrt{7}\] \[\Large 2\sqrt{7}\] So \[\Large \sqrt{28}=2\sqrt{7}\]

OpenStudy (anonymous):

so i wright all that

jimthompson5910 (jim_thompson5910):

If you want. Those are just the steps to getting to that answer.

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

how do i type the divistion thing

jimthompson5910 (jim_thompson5910):

if you want to say something like "2 divided by 3", then you would type "2/3"

OpenStudy (anonymous):

k